# A Survey on Large Language Model based Autonomous Agents

### v4

@@ 2.1.4 Action Module
Operation
Structure
WebGPT [66]
12/2021
SayCan [78]
04/2022
MRKL [72]
05/2022
Inner Monologue [61]
07/2022
Social Simulacra [79]
08/2022
ReAct [59]
10/2022
MALLM [43]
01/2023
DEPS [33]
02/2023
Toolformer [15]
02/2023
Reflexion [12]
03/2023
CAMEL [80]
① ②
03/2023
API-Bank [69]
04/2023
ViperGPT [74]
03/2023
HuggingGPT [13]
03/2023
Generative Agents [20]
04/2023
LLM+P [57]
04/2023
ChemCrow [75]
04/2023
OpenAGI [73]
04/2023
AutoGPT [81]
04/2023
SCM [35]
04/2023
Socially Alignment [82]
05/2023
GITM [16]
05/2023
Voyager [38]
05/2023
Introspective Tips [83]
05/2023
RET-LLM [42]
05/2023
ChatDB [40]
06/2023
$S^{3}$ [77]
07/2023
ChatDev [18]
07/2023
ToolLLM [14]
07/2023
MemoryBank [39]
07/2023
MetaGPT [23]
08/2023
@@ 3.1 Social Science
[…]
@@ 3.2 Natural Science
Domain
Work
Psychology
TE [102], Akata et al. [103], Ziems et al. [105], Ma et al. [104]
Social Science
Political Science and Economy
Out of One [29], Horton [106], Ziems et al. [105]
Social Simulation
[…]
Jurisprudence
ChatLaw [112], Blind Judgement [113]
Research Assistant
Ziems et al [105], Bail et al. [114]
Documentation and Data Management
ChemCrow [75], Boiko et al. [115]
Natural Science
Experiment Assistant
ChemCrow [75], Boiko et al. [115], Grossmann et al. [122]
Natural Science Education
ChemCrow [75], CodeHelp [120], Boiko et al. [115], MathAgent [117], Drori et al. [118]
Engineering
CS & SE
RestGPT [70], Self-collaboration [24], SQL-PALM [90], RAH [92], DB-GPT [41], RecMind [51], ChatEDA [123], InteRecAgent [124], PentestGPT [125], CodeHelp [120], SmolModels [126], DemoGPT [127], GPTEngineer [128]
Industrial Automation
GPT4IA [129], IELLM [130], TaskMatrix.AI [71]
Robotics & Embodied AI
ProAgent [131], LLM4RL [132], PET [133], REMEMBERER [134], DEPS [33], Unified Agent [135], SayCan [78], LMMWM [136], TidyBot [137], RoCo [93], SayPlan [31]
@@ 4.1 Subjective Evaluation
Model
Subjective
Objective
Benchmark
Time
WebShop [85]
① ③
$\checkmark$
07/2022
Social Simulacra [79]
08/2022
TE [102]
08/2022
LIBRO [162]
09/2022
ReAct [59]
$\checkmark$
10/2022
Out of One, Many [29]
② ③
02/2023
DEPS [33]
$\checkmark$
02/2023
Jalil et al. [163]
02/2023
Reflexion [12]
① ③
03/2023
IGLU [138]
$\checkmark$
04/2023
Generative Agents [20]
① ②
04/2023
ToolBench [153]
$\checkmark$
04/2023
GITM [16]
$\checkmark$
05/2023
Two-Failures [164]
05/2023
Voyager [38]
$\checkmark$
05/2023
SocKET [165]
② ③
$\checkmark$
05/2023
MobileEnv [166]
① ③
$\checkmark$
05/2023
Clembench [167]
① ③
$\checkmark$
05/2023
Dialop [168]
$\checkmark$
06/2023
Feldt et al. [169]
06/2023
CO-LLM [22]
07/2023
Tachikuma [170]
$\checkmark$
07/2023
WebArena [171]
$\checkmark$
07/2023
RocoBench [93]
① ② ③
07/2023
AgentSims [34]
08/2023
AgentBench [172]
$\checkmark$
08/2023
BOLAA [173]
① ③ ④
$\checkmark$
08/2023
Gentopia [174]
$\checkmark$
08/2023
EmotionBench [160]
$\checkmark$
08/2023
PTB [125]
08/2023

### v5

@@ 2.1.4 Action Module
Operation Structure
WebGPT [66] - - - - 12/2021
SayCan [78] - - - 04/2022
MRKL [72] - - - - 05/2022
Inner Monologue [61] - - - 07/2022
Social Simulacra [79] - - - - 08/2022
ReAct [59] - - - 10/2022
MALLM [43] - - 01/2023
DEPS [33] - - - 02/2023
Toolformer [15] - - - 02/2023
Reflexion [12] - 03/2023
CAMEL [80] ① ② - - - 03/2023
API-Bank [69] - - - 04/2023
ViperGPT [74] - - - - - 03/2023
HuggingGPT [13] - - - 03/2023
Generative Agents [20] - 04/2023
LLM+P [57] - - - - 04/2023
ChemCrow [75] - - - - 04/2023
OpenAGI [73] - - - 04/2023
AutoGPT [81] - 04/2023
SCM [35] - - - 04/2023
Socially Alignment [82] - - 05/2023
GITM [16] - 05/2023
Voyager [38] - 05/2023
Introspective Tips [83] - - - 05/2023
RET-LLM [42] - - 05/2023
ChatDB [40] - - 06/2023
$S^{3}$ [77] - - 07/2023
ChatDev [18] 07/2023
ToolLLM [14] - - - 07/2023
MemoryBank [39] - - - 07/2023
MetaGPT [23] - 08/2023
@@ 3.1 Social Science
Social Simulation: Previously, conducting experiments with human societies is often expensive, unethical, or even infeasible. With the ever prospering of LLMs, many people explore to build virtual environment with LLM-based agents to simulate social phenomena, such as the propagation of harmful information, and so on [79, 107, 108, 20, 34, 109, 77, 110]. For example, Social Simulacra [79] simulates an online social community and explores the potential of utilizing agent-based simulations to aid decision-makers to improve community regulations.  [107, 108] investigates the potential impacts of different behavioral characteristics of LLM-based agents in social networks. Generative Agents [20] and AgentSims[34] construct multiple agents in a virtual town to simulate the human daily life. SocialAI School [109] employs LLM-based agents to simulate and investigate the fundamental social cognitive skills during the course of child development. S3 [77] builds a social network simulator, focusing on the propagation of information, emotion and attitude. CGMI [111] is a framework for multi-agent simulation. CGMI maintains the personality of the agents through a tree structure and constructs a cognitive model. The authors simulated a classroom scenario using CGMI.
@@ 3.2 Natural Science
Domain Work
Psychology TE [102], Akata et al. [103], Ziems et al. [105], Ma et al. [104]
Social Science Political Science and Economy Out of One [29], Horton [106], Ziems et al. [105]
Social Simulation Social Simulacra [79], Generative Agents [20], SocialAI School [109], AgentSims [34], S3 [77], Williams et al. [110], Li et al. [107], Chao et al. [108]
Jurisprudence ChatLaw [112], Blind Judgement [113]
Research Assistant Ziems et al [105], Bail et al. [114]
Documentation and Data Management ChemCrow [75], Boiko et al. [115]
Natural Science Experiment Assistant ChemCrow [75], Boiko et al. [115], Grossmann et al. [122]
Natural Science Education ChemCrow [75], CodeHelp [120], Boiko et al. [115], MathAgent [117], Drori et al. [118]
Engineering CS & SE RestGPT [70], Self-collaboration [24], SQL-PALM [90], RAH [92], DB-GPT [41], RecMind [51], ChatEDA [123], InteRecAgent [124], PentestGPT [125], CodeHelp [120], SmolModels [126], DemoGPT [127], GPTEngineer [128]
Industrial Automation GPT4IA [129], IELLM [130], TaskMatrix.AI [71]
Robotics & Embodied AI ProAgent [131], LLM4RL [132], PET [133], REMEMBERER [134], DEPS [33], Unified Agent [135], SayCan [78], LMMWM [136], TidyBot [137], RoCo [93], SayPlan [31]
@@ 4.1 Subjective Evaluation
Model Subjective Objective Benchmark Time
WebShop [85] - ① ③ $\checkmark$ 07/2022
Social Simulacra [79] - 08/2022
TE [102] - - 08/2022
LIBRO [162] - - 09/2022
ReAct [59] - $\checkmark$ 10/2022
Out of One, Many [29] ② ③ - 02/2023
DEPS [33] - $\checkmark$ 02/2023
Jalil et al. [163] - - 02/2023
Reflexion [12] - ① ③ - 03/2023
IGLU [138] - $\checkmark$ 04/2023
Generative Agents [20] ① ② - - 04/2023
ToolBench [153] - $\checkmark$ 04/2023
GITM [16] - $\checkmark$ 05/2023
Two-Failures [164] - - 05/2023
Voyager [38] - $\checkmark$ 05/2023
SocKET [165] - ② ③ $\checkmark$ 05/2023
MobileEnv [166] - ① ③ $\checkmark$ 05/2023
Clembench [167] - ① ③ $\checkmark$ 05/2023
Dialop [168] - $\checkmark$ 06/2023
Feldt et al. [169] - - 06/2023
CO-LLM [22] - 07/2023
Tachikuma [170] $\checkmark$ 07/2023
WebArena [171] - $\checkmark$ 07/2023
RocoBench [93] - ① ② ③ - 07/2023
AgentSims [34] - - 08/2023
AgentBench [172] - $\checkmark$ 08/2023
BOLAA [173] - ① ③ ④ $\checkmark$ 08/2023
Gentopia [174] - $\checkmark$ 08/2023
EmotionBench [160] - $\checkmark$ 08/2023
PTB [125] - - 08/2023